OpenStudy (anonymous):

HELP PLEASE!! Find the cube roots of 8(cos 216° + i sin 216°).

OpenStudy (anonymous):

take the cube root of 8, that is 2 then divide each angle by 3

OpenStudy (anonymous):

that is why this is easy, why you want to write a complex number in trig form \[216\div 3=72\] so one answer is \[2\left(\cos(72)+i\sin(72)\right)\]

OpenStudy (michele_laino):

please you have to apply the De Moivre formula, namely: \[\sqrt[3]{8}(\cos (\frac{ 216+2k 180 }{ 3 })+i \sin(\frac{ 216 +2k180}{ 3 }))\] and substitute k=0,1,2 so you should get three square roots

OpenStudy (anonymous):

to find the other one, go around the circle again i.e. add 360 to 216, then divide the result by 3 lather rinse repeat

OpenStudy (anonymous):

[8^(1/3) [\cos1/3(216\deg+360k) + i \sin 1/3(216\deg+360k)]

OpenStudy (anonymous):

Right?

OpenStudy (michele_laino):

right!

OpenStudy (michele_laino):

for example, for k=0, you should get this: \[2(\cos(72)+i \sin(72))\]

OpenStudy (anonymous):

So, = 2[cos (72deg+ 120degk) + i sin (72deg + 120degk)] Right?

OpenStudy (anonymous):

So, k = 0: 2(cos 72deg + i sin 72deg) right?

OpenStudy (michele_laino):

no, please if I set k=1 in my formula, I get this: \[2\left( \cos \left( \frac{ 216+1*360 }{ 3 } \right)+i \sin \left( \frac{ 216+1*360 }{ 3 } \right) \right)\]

OpenStudy (anonymous):

oh ok

OpenStudy (michele_laino):

and if I set k=2, I get: \[2\left( \cos \left( \frac{ 216+2*360 }{ 3 } \right)+ i \sin \left( \frac{ 216+2*360 }{ 3 } \right) \right)\]

OpenStudy (anonymous):

So, k = 2: 2(cos 312deg + i sin 312deg)

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

yes!

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

thank you!

OpenStudy (michele_laino):

thank you!